The area of a circle is 28 square meters. What is its circumference?
Answer by stanbon(75887)

The area of a circle is 28 square meters. What is its circumference?
------
Area = (pi)r^2
28 = (pi)r^2
r = sqrt[28/pi] meters
--------------------
Circumference = 2(pi)r
C = 2(pi)*sqrt(28/pi)
C = 2*sqrt(28)
C = 2*2sqrt(7)
C = 4*sqrt(7) meters
--------------------
Cheers,
Stan H.